# Synthesising Activity Participations and Scheduling with Deep Generative Machine Learning

### v2

@@ 8 Sample Size Scenarios
Table 15: Continuous RNN Sub-Model Hyper-Parameters
@@
Table 16: Sample Size Scenarios Evaluation Summary
@@
Figure 16: Sample Sizes Scenarios: Aggregate Activity Frequencies (10 minute time bins)
Figure 17: Sample Sizes Scenarios: Activity Sequences
9 COVID-19 Scenario
Finally, we consider applying the Continuous RNN model to a behavioural change scenario. We extract schedules from the NTS travel diaries as per Section 2.5, but create two new real samples for 2019 and 2020. We chose these samples as we expect a change in schedule distributions due to COVID-19 from 2019 to 2020. We name these samples Real 2019 and Real 2020. Both samples have 23,452 schedules, significantly less than the base 2023 sample, so we train a Continuous RNN Mid (Table 15) for each sample year, then generate synthetic samples from both. We name these samples Synthetic 2019 and Synthetic 2020.
@@ 9 COVID-19 Scenario
Both synthetic samples (2019 and 2020) show lower levels of participation than their respective real samples. However, the change from 2019 to 2020 is well represented in the synthetic populations. Most notably, from 2019 to 2020, the real expected sequence length drops from 3.887 to 3.717 (-0.171). Although the synthetic samples both under-estimate sequence length, they show a similar drop from 3.700 to 3.527 (-0.173). Single and pair participations show a similar pattern, suggesting that although the models appear biased to lower activity participation, this bias is consistent across scenarios, such that the behaviour change is captured.
Transitions are extremely similar between the real samples, so they are not considered. For timing estimation, the models capture both real sample distributions well and therefore also show similar changes between 2019 and 2020. For example, there is a shift to activities starting 0.016 days (23 minutes) earlier in real samples, and 0.015 days (22 minutes) earlier in the synthetic samples.
This experiment suggests that the approach can capture differences between samples of schedules. Specifically, we show that the model design has not been overfit to the original 2023 scenario and can be re-trained to capture new patterns of schedules, such as for 2019 and 2020. However, we also show that the reported differences between synthetic and real samples, particularly for participations, are comparable to changes observed between 2019 and 2020 due to COVID-19.
Table 17: COVID Scenario Mean Densities
Real Synthetic* Unit
@@
* mean of mean density from 5 model runs
@@ 10 Results Discussion
Table 18: Model Sizes and Run Times
@@
Experiments with smaller data samples in Section 8 show that more data leads to better results. Smaller models can be used for smaller datasets, but the deep generative approach becomes unviable at around 7,500 schedules or fewer. Our COVID-19 scenario in Section 9 shows that the Continuous RNN model design and hyperparameter specification can capture differences between samples with different patterns of schedules. However, this scenario also suggests that the expected differences in participation rates may be non-trivial.
@@ 11 Conclusions
We demonstrate the synthesis of large, diverse samples of realistic activity schedules with a deep generative approach based on a VAE architecture. Our work shows that a deep generative modelling approach can approximate a complex human behavioural system. Our generative approach makes rapid and diverse synthesis, upsampling or anonymisation of observed activity schedules feasible. Such samples can then be used for applied models and simulations in the transport, energy and epidemiology domains.

### v3

@@ 8.1 Sample Size Scenarios
Table 18: Sample Size Scenarios Hyperparameters: ContRNN Mid, Small & Tiny
@@
Table 19: Sample Size Scenarios Evaluation Summary: Domain-level distances such that lower is better
@@
Figure 15: Sample Sizes Scenarios: Aggregate activity frequencies
Figure 16: Sample Sizes Scenarios: Disaggregate sequences of activity types
8.2 Alternative Year Scenarios
We evaluate how well the Continuous RNN performs on new datasets it hasn’t seen before. Specifically, we train and test the model using data from 2019 and 2020. For each dataset, we retrain the model without changing any of its hyperparameters. This approach shows that the model architecture and hyperparameters are not overfit to the original data and that it can generalise effectively to new situations.
We extract schedules from the NTS travel diaries as per Section 2.5, but create two new real samples for 2019 and 2020. We chose these samples as we expect a change in schedule distributions due to COVID-19 from 2019 to 2020. We name these samples Real 2019 and Real 2020. Both samples have 23,452 schedules, significantly less than the base 2023 sample, so we train a Continuous RNN Mid (Table 18) for each sample year, then generate synthetic samples from both. We name these samples Synthetic 2019 and Synthetic 2020.
@@ 8.2 Alternative Year Scenarios
Both synthetic samples (2019 and 2020) show lower levels of participation than their respective real samples. This is consistent with previous results in Section 7.1 and can likely be improved with more training data as per Section 8.1. The change from 2019 to 2020 is well represented in the synthetic populations. Most notably, from 2019 to 2020, the real expected sequence length drops from 3.887 to 3.717 (-0.171). The synthetic samples show a similar drop from 3.700 to 3.527 (-0.173). Single and pair participations show a similar pattern, suggesting that although the models appear biased to lower activity participation, this bias is consistent across scenarios, such that the model responds to new data and therefore behaviour change is captured.
Transitions are extremely similar between the real samples, so they are not considered. For timing estimation, the models capture the real sample distributions well and therefore also show similar changes between 2019 and 2020. For example, there is a shift to activities starting 0.016 days (23 minutes) earlier in real samples, and 0.015 days (22 minutes) earlier in the synthetic samples.
This experiment suggests that the approach can capture differences between samples of schedules. Specifically, we show that the model design (both architecture and hyperparameters) has not been overfit to the original 2023 scenario and can be re-trained to capture new patterns of schedules, such as for 2019 and 2020. However, we also show that the reported differences between synthetic and real samples, particularly for participations, although small, are comparable to changes observed between 2019 and 2020 due to COVID-19.
Table 20: Alternative Year Scenarios: 2019 & 2020 distribution-level descriptions (mean densities) and differences
Real Synthetic Unit
@@
(All results are mean descriptive metrics from 5 model runs)
8.3 Comparison with Supervised Discriminative Approach
Table 21: UK National Travel Survey Label Summary
Attribute Categories
Gender {male, female, unknown}
Age {0-4, 5-10, 11-15, 16-19, 20-29, 30-39, 40-49, 50-69, 70+, unknown}
Car Access {yes, no, unknown}
Work Status {employed, education, unemployed}
Income (household) {highest, high, medium, low, lowest, unknown}
Higher Education {yes, no, unknown}
License {yes, no, unknown}
Area {suburban, urban, rural, Scotland, unknown}
Ethnicity {white, non-white, unknown}
Household Composition {1A, 2A, 3A+, 1A&1C+, 2A&1C+, 3A+&1C+, unknown}*
WfH {all, most, some, none, unknown}
Mobility Difficulty {none, all, walk&bus, walk&car, bus&car, walk, bus, car, unknown}
Wheelchair User {yes, no, unknown }
Season Ticket {subsidised, seasonal, other, none, unknown}
Health {very good, good, fair, bad, very bad, unknown}
Blue Badge {yes, no, unknown}
* A: adult, C: child
Figure 17: Comparison of Real, ContRNN and Discriminative Models: Aggregate activity frequencies
Figure 18: Comparison of Real, ContRNN and Discriminative Models: Disggregate sequences of activity types
To demonstrate the advantages of our generative approach for synthesising realistic distributions of schedules over discriminative approaches, we compare our best-performing generative model, ContRNN, against two supervised discriminative models [^5] We note that no code has been made available for the previous ML activity scheduling approaches discussed in Section 1.7. We have therefore replicated the methodology of Koushik et al. (2023) who describe the experimental methodology in sufficient detail to be able to do so.: (i) a discriminative model developed using the methodology of Koushik et al. (2023) with discretised time, which we call Discrim, and (ii) a second discriminative model that makes use of the continuous activity encoding, which we call ContDiscrim.
Both the discriminative models predict schedules based on input sociodemographic attributes. Specifically, they seek to infer the most likely activity (and duration) at each step in a sequence, based on input attributes. Both models are trained using a standard supervised training dataset of paired attributes and schedules. To create a training dataset, we pair the NTS schedules, described in Section 2.5, with their associated attributes, extracted from NTS individual and household data tables, and described in Table 21. Attributes are treated as nominal categories, one-hot encoded and then concatenated to create an input vector for each schedule.
The Discrim model is described in detail by Koushik et al. (2023). Notably, schedules use the discrete encoding with 10-minute steps. The model is composed of three bidirectional LSTM layers of size 200 and 30% dropout. Activity predictions at each step are made as per the discrete embedding described in Section 4.2.1. The model is trained to minimise mean cross-entropy. Individual activity losses are weighted by inverse frequency in the training dataset.
The ContDiscrim model updates the design of the Discrim model to a continuous schedule encoding. This performs better than the Discrim model and enables fairer comparison to the ContRNN VAE. The model is composed of four LSTM layers of size 128 and 10% dropout. Activity and duration predictions are made as per the continuous embedding in Section 3.2. The loss function includes duration loss as per the continuous models reconstruction loss in Section 4.3.
For evaluation, we train the models, generate schedules, and then evaluate them using our generative evaluation methodology described in Section 5. To simulate the generative or synthesis case, the discriminative models generate schedules from the training dataset attributes. Note that in a typical supervised learning framework, we would seek to evaluate against a withheld dataset of attributes and schedules. However, in our experiment, we consider a generative or synthesis application. This also allows comparison with the generative ContRNN VAE.
Table 22: Comparison of ContRNN to discriminative models: domain-level distances such that lower is better
Generative Discriminative
ContRNN ContDiscrim Discrim Unit
Density Estimation
Participations 0.062 0.377 0.796 rate EMD
Transitions 0.004 0.018 0.026 rate EMD
Timing 0.024 0.152 0.196 days EMD
Sample Quality
Invalid 0.019 0.002 0.004 prob.
Creativity
Homogenous 0.022 0.879 0.975 prob.
Conservative 0.013 0.016 0.402 prob.
(All results are mean distances from 5 model runs)
Figures 18 and 18 show that both discriminative models fail to create a realistic distribution of schedules. This is confirmed by domain-level density evaluation in Table 22. We can explain the poor performance of the Discrim model as a combination of the discrete embedding and relatively long RNN, which bias the model towards predicting the most likely activity at each time step, which is home. This is discussed in more detail in Section 6.1.
The ContDiscrim model also performs worse than the ContRNN, despite being architecturally comparable, suggesting that without the explicit incorporation of a generative capability (through the latent space in the VAE), the models struggle to learn a realistic distribution of schedules.
@@ 9 Results Discussion
Table 23: Model Sizes and Run Times for different generative model architectures
@@
Experiments with smaller data samples in Section 8.1 show that more data leads to better results. Smaller models can be used for smaller datasets, but the deep generative approach becomes unviable at around 7,500 schedules or fewer. Our scenarios in Section 8.2 show that the Continuous RNN model architecture and hyperparameters can reasonably be expected to perform well on datasets in future years.
Section 8.3 shows that the generative approach is more capable of modelling the variance of real schedules than existing discriminative approaches. This is unsurprising given the complexity of the real scheduling process and the relatively limited availability of explanatory variables. We effectively demonstrate that variance in activity schedules cannot be well modelled using available attributes alone, that discriminative models would likely benefit from incorporating generative mechanisms, and that the ContRNN should likely be favoured over discriminative models when seeking to synthesise a realistic distribution of schedules from data.
@@ 10 Conclusions
We demonstrate the synthesis of large, diverse samples of realistic activity schedules with a deep generative approach based on a VAE architecture. Our work demonstrates a deep generative ML model approximating a human behavioural system, with complex dependencies and constraints. Our generative approach makes rapid and diverse synthesis, upsampling or anonymisation of observed activity schedules feasible. Such samples can then be used for applied models and simulations in the transport, energy and epidemiology domains.